There are three triples of positive integers (a,b,c, d,e,f, and g,h,i) such that a^2+b^2+c^2 = d^2+e^2+f^2 = g^2+h^2+i^2 = 101. Evaluate the expression abc+def+ghi.
I've found three
1. 101 = 1^2 + 6^2 + 8^2
2. 101 = 2^2 + 4^2 + 9^2
3. 101 = 4^2 + 6^2 + 7^2
They all share terms so I don't think these are options.

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52781)

.

Three triples are

(a,b,c) = (1,6,8) :   1^2 + 6^2 + 8^2 = 101


(d,e,f) = (2,4,9) :   2^2 + 4^2 + 9^2 = 101


(g,h,i) = (4,6,7) :   4^2 + 6^2 + 7^2 = 101


The sum  abc + def + ghi = 1*6*8 + 2*4*9 + 4*6*7 = 288.    ANSWER

Solved.



Answer by Edwin McCravy(20055)

Triples that share 1 or 2 numbers are still DIFFERENT triples. Triples would have
to share ALL THREE numbers to be the same triple.

Edwin